OpenStudy (anonymous):

HELP!! simplify 9^1/3 x 81^1/3

OpenStudy (nubeer):

hmm you can start trying like this \[9^{\frac{ 1 }{ 3 }}= 3^{2\times \frac{ 1 }{ 3 }} = 3^{\frac{ 2 }{ 3 }}\] you have to do same with 81

OpenStudy (anonymous):

then what do i do?

OpenStudy (nubeer):

after u find out.. then you would have same base 3. so u can add the power of 3

OpenStudy (nubeer):

\[3^{2}\times 3^{3} = 3^{5}\]

OpenStudy (anonymous):

alright got it